OpenStudy (fanduekisses):

Did I do this right? y=2(x^2-3x) Find dy/dt when x=3 given that dx/dt=2

OpenStudy (fanduekisses):

dy/dt=2*2x*dx/dt * 3 dy/dt=12x * dx/dt dy/dt=72

OpenStudy (anonymous):

dy/dt = dy/dx * dx/dt

OpenStudy (fanduekisses):

I used the product rule

OpenStudy (anonymous):

y = 2x^2-6x dy/dt = 4x*dx/dt-6*dx/dt since dx/dt = 2 and x = 3 => dy/dt = 4(2x-3) = 4(6-3) = 4*3 = 12